Good afternoon ^.^ i was wondering if someone could show me how to solve this problem : "when 75 is subtracted from four times Tamu's age, the result is Tamu's age. How old is Tamu?" i'd appreciate it :) please and thank you!
Answer by jim_thompson5910(35256)

Let T = Tamu's age

"when 75 is subtracted from four times Tamu's age, the result is Tamu's age"

translates to

4T - 75 = T

Let's solve for T

4T - 75 = T

4T - 75 - 4T = T - 4T ... Subtract 4T from both sides.

-75 = -3T ... Combine like terms.

-3T = -75 ... Flip the equation

-3T/(-3) = -75/(-3) ... Divide both sides by -3.

T = 25 ... Divide

So Tamu is 25 years old

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------